# Amortized Sampling with Transferable Normalizing Flows

### v2

@@ 4.6 Inference-time temperature transfer
Figure 5: Scaled prior greatly the ability of Prose to accurately reweight to arbitrary temperatures. Metrics for Prose on RLMM unseen tetrapeptide, targeting temperatures up to $800\text{\,}\mathrm{K}$. Naively applying SNIS to the target temperature leads to a rapid degradation in energy distribution, and to a lesser extent the dihedral angle distribution. Applying prior scaling (Prose SP) leads to a significant improvement in energy distribution at high temperatures and moderate improvement in dihedral angles. Notably, the TICA distribution improves at higher temperatures irrespective of scaled prior usage, although scaled prior remains more effective.
@@ 5 Related Work
Boltzmann generators are machine learning-based samplers that train likelihood-based models and employ inference-time SNIS to achieve consistent sampling of the target density (Noé et al., 2019). A major limitation of standard Boltzmann generators is the need to train the proposal model on a dataset of true density samples, motivating methods that transfer between molecular systems. Accordingly, Jing et al. (2022) propose a transferable Boltzmann generator operating on the torsion angles of small molecules, and Klein and Noe (2024) develop a Boltzmann generator operating on Cartesian coordinates that transfers between dipeptide systems. However, the scalability of such methods has remained limited due to the difficulty of designing expressive generative models that possess efficient and accurate likelihood evaluations. In particular the use of continuous normalizing flows implies a large cost to proposal likliehoods due to the need to integrate the vector field divergence Grathwohl et al. (2019). Schopmans and Friederich (2025) replace the single-step SNIS of standard Boltzmann generators with a temperature-annealing sequence of normalizing flows, performing SNIS with each flow to sample from a given target distribution.
@@ Acknowledgments
Appendix A Dataset
@@
Local energy minimization is performed with the Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BGFS) algorithm. Energy minimization is followed by burn-in simulation of length $50\text{\,}\mathrm{ps}$, after which samples are collected every $1\text{\,}\mathrm{ps}$ (train) or $10\text{\,}\mathrm{ps}$ (test) until the simulation budget is exhausted. Full MD simulation parameters are provided in Table˜5.
@@
Appendix B Prose Architecture Details
Tokenization and masking
[…]
with log-determinant of Jacobian given by
$\log\mathopen{}\mathclose{{\left|\frac{\partial f_{t}(z_{t})}{\partial z_{t}}}}\right|=-\sum_{i=1}^{d-1}m[i]\cdot\alpha_{t}(z_{t}[:i])[i].$ (12)
[…]
Permutations
All permutations are defined from the N-terminal to the C-terminal. In the residue-by-residue permutation the atoms are ordered such that each residue forms a contiguous sequence, with sidechain atoms immediately following the corresponding backbone atoms. In the backbone-first permutation the entire sequence of backbone atoms is placed at the start of the sequence before any sidechain atoms. Where constituent residues possess a branch or ring we also introduce a variant in which the branch ordering is flipped or ring traversal is inverted. The flip permutations are a simple inversion of the permutation, where padding tokens and not moved from their position at the end of the sequence. The specific sequence of permutations employed in Prose is presented in Table˜7.
Permutation Description
$\pi_{0}$ Backbone first
$\pi_{1}$ Residue–by–residue (flip)
$\pi_{2}$ Backbone–first (flip)
$\pi_{3}$ Residue–by–residue
$\pi_{4}$ Backbone–first (variant)
$\pi_{5}$ Residue–by–residue (variant, flip)
$\pi_{6}$ Backbone–first (variant, flip)
$\pi_{7}$ Residue–by–residue (variant)
Adaptive Layer Norm and Transition
Appendix C Importance sampling variants
C.1 Self-normalized importance sampling (SNIS)
SNIS corresponds to the following estimator
@@ C.1 Self-normalized importance sampling (SNIS)
C.2 Continuous-time annealed importance sampling (AIS)
@@
C.3 Discrete-time annealed importance sampling (AIS)
@@
C.4 Sequential Monte Carlo
@@
Appendix D Training configuration
All models are trained for $5\cdot 10^{5}$ iterations using a batch size of 512 with the AdamW optimizer [Loshchilov and Hutter, 2018]. We employ a cosine learning rate schedule in which the initial and final learning rates are a reduction of the maximal value by factor of 500, as well as exponential moving average with decay of 0.999. No overfitting was observed hence no early stopping was required. Samples are normalized using the standard deviation of the longest sequence data in a given experiment, noting that a single value must be shared across systems of different dimensionality. An overview of all training configurations is provided in Table˜8.
Continuous Normalizing Flows
TarFlows
Following Zhai et al. [2024] and Tan et al. [2025] we use a learning rate of $1\cdot 10^{-4}$, weight decay of $4\cdot 10^{-4}$, and AdamW $\beta_{1},\beta_{2}$ of $(0.9,0.95)$. Data augmentation is applied as random rotations and Gaussian center of mass augmentation, in which every the entire system conformation is translated by a vector $c\sim\mathcal{N}(0,\sigma^{2}I_{3})$. The $\sigma^{2}$ value is chosen to match that of the prior, which has a center of mass $c^{2}=\frac{1}{N}$ where $N$ is the number of atoms. Given $N$ is in our case variable for a single model trained on multiple systems, this augmentation is applied-per system before padding is applied. The architecture width and depth is provided in Table˜9.
Table 8: Overview of training configurations.
ECNF ECNF++ TarFlow / Prose
Learning Rate $5\cdot 10^{-4}$ $5\cdot 10^{-4}$ $1\cdot 10^{-4}$
Weight Decay 0.0 $1\cdot 10^{-2}$ $4\cdot 10^{-4}$
$\beta_{1},\beta_{2}$ 0.9, 0.999 0.9, 0.999 0.9, 0.95
EMA Decay 0.0 0.999 0.999
ECNF ECNF++ TarFlow Prose
Channels 128 256 384 384
Depth 9 9 8 8
Layers per block N/A N/A 8 8
Parameters (M) 1 4 115 285
D.1 Computational resources
ECNF++ TarFlow Prose
Training iterations / H100 hour 960 1132 260
@@ E.1 Proposal sampling and likelihood evaluation
$\log p^{c}_{\theta}(x)=\log p_{\theta}(x)+\frac{\|c\|^{2}}{2\sigma^{2}}-\log\mathopen{}\mathclose{{\left[\frac{\|c^{2}\|}{\sqrt{2}\sigma^{3}\Gamma\mathopen{}\mathclose{{\left(\frac{3}{2}}}\right)}}}\right],$ (37)
where $\Gamma$ is the gamma function. This adjustment seeks to account for the radial component of the proposal density. We additionally apply the same weight clipping threshold as in the ECNF when performing SNIS, or before SMC.
@@ E.2 Metrics
We compare generated samples to ground truth data, collected as defined in Appendix˜A, using empirical Wasserstein-2 distances. Given empirical distributions $\mu=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ and $\nu=\frac{1}{m}\sum_{j=1}^{m}\delta_{y_{j}}$, the empirical Wasserstein-2 distance is defined as
@@
Following Raja et al. [2025], on the reference MD trajectory we run $k$-means ($k=20$) on the first two time-lagged independent components (TICs) fitted to the pairwise distances and dihedral angles of the peptide confirmations. Using these clusters, we obtain the occupancy distributions for the reference MD and the samples generated from Prose, and report their Jenson-Shannon divergence ($\mathrm{TICA}\text{‑}\mathrm{JSD}$), defined as
@@
where $P$ and $Q$ are discrete distributions, and $M=\frac{1}{2}(P+Q)$. Moreover, we repeat the same procedure using $k$-means only on the dihedral angles, and report this as $\mathcal{T}\text{‑}\mathrm{JSD}$. In both cases we fit the $k$-means clustering on the features (TICA projections or dihedrals) from full reference MD, but compute the metrics of generated samples against a subsampling of the reference trajectory as previously discussed. Note that JSD depends on the arbitrary choice of clustering (here $k=20$) and can be sensitive to binning resolution, whereas Wasserstein distances avoid discretization by operating directly on the continuous distributions.
@@ E.3 Sampling algorithm configurations
For both variants we generate a proposal set of $10^{4}$ samples. For the discrete variant, we perform 50 annealing steps, requiring two energy evaluations per step: one to update the samples and one for the Metropolis–Hastings correction. For the continuous variants, we perform 100 annealing steps. In both cases, resampling is performed at every step. For the continuous variant, Langevin dynamics is used with a step size $\sigma_{t}$ of $10^{-7}$ for dipeptides and tetrapeptides, and $\sigma_{t}$ of $10^{-8}$ for octopeptides. Details of the formulation are found in Section˜C.2. For the discrete variant, we apply Langevin dynamics with a step size of $10^{-5}$, followed by a Metropolis-Hastings step to accept or reject proposals. The step size is adaptively updated to maintain an acceptance rate of approximately 60%, under the assumption of sufficient smoothness in the intermediate densities. Further details of discrete-time AIS are found in Section˜C.3.
@@ E.4 Additional baseline configurations
The inference code of Lewis et al. [2024] was employed. BioEmu does not directly mode all-atom resolution hence hpacker is employed to introduce the side chains before energy minimization. The codebase of [Lewis et al., 2024] was adapted to use the same Amber14 forcefield as ManyPeptidesMD, on which the $10^{4}$ energy evaluation budget was spent on per-sample minimization. We additionally experimented with equilibration to target the larger $10^{6}$ energy budget but could not achieve superior results to minimization-only hence this was not included. The adapted codebase for this baseline is provided at https://github.com/transferable-samplers/BioEmu.
@@ E.5 Computational resources
Appendix F Additional results
F.1 Proposal Performance
@@ F.2 Additional ablations
We consider three further ablations of the Prose architecture. Firstly, we replace the adaptive conditioning and transition blocks with (i) deeper transformation blocks in which the 8 transformer layers are increased to 20 (ii) a wider transformer block in which the dimension is increased from 384 to 576. In both cases the increased depth / width was defined to approximately match the parameter count of Prose. We additionally ablate the use of “look-ahead conditioning”, in which the atom token $z[i]$ is conditioned not only on $A_{i},R_{i},P_{i},L_{i}]$ but also $A_{i+1},R_{i+1},P_{i+1},L_{i+1}]$, with conditioning information for this pair of indexes mixed using a small MLP. This was motivated by the observation that given naive conditioning the causal masking implies updates to $z[i]$ are computed without knowledge of $A_{i},R_{i},P_{i},L_{i}]$. Preliminary experiments up to tetrapeptides suggested look-ahead conditioning to be beneficial hence it was included in the final Prose model.
@@ F.3 Dataset scaling
We explore the effect of dataset size on the performance of Prose. The full ManyPeptidesMD dataset contains 21,700 sequences simulated for 200ns each. We train a further four models wherein (i) the trajectories are limited to 50ns and 100ns respectively, and (ii) the set of sequences is reduced to 25% and 50%. For the 25% and 50% sequence reduction variants sequence reduction is applied non-linearly with a simple estimate of relative cost, to avoid excessively removing the relatively cheap simulations of short sequences. In Fig.˜9 we present metrics for these models. We note the $E\text{‑}\mathcal{W}_{2}$ to lack any meaningful trend, and the $\mathcal{T}\text{‑}\mathcal{W}_{2}$ to have limited sensitivity to the dataset variation, despite a slight trend towards larger data subsets improving the metrics. Most interesting of these plots is the $\mathrm{TICA}\text{‑}\mathcal{W}_{2}$ with a clear trend indicating the full 200ns trajectories to be beneficial, whilst using only 50% of the sequences is in fact superior to the full dataset.
Figure 9: Wasserstein-distance metrics for Prose trained on variants of ManyPeptidesMD. Upper row: ManyPeptidesMD contains 200ns trajectories for all training sequences, we train on a variant with 50ns and 200ns respectively. Lower: ManyPeptidesMD contains 21,700 sequences, we train using 25% and 50% of the total. Evaluation metrics computed using 10 test octopeptide sequences.
F.4 JSD evaluation

### v3

@@ 4.6 Inference-time temperature transfer
Figure 5: Scaled prior greatly improves the ability of Prose to accurately reweight to arbitrary temperatures. Metrics for Prose on RLMM unseen tetrapeptide, targeting temperatures up to $800\text{\,}\mathrm{K}$. Naively applying SNIS to the target temperature leads to a rapid degradation in energy distribution, and to a lesser extent the dihedral angle distribution. Applying prior scaling (Prose SP) leads to a significant improvement in energy distribution at high temperatures and moderate improvement in dihedral angles. Notably, the TICA distribution improves at higher temperatures irrespective of scaled prior usage, although scaled prior remains more effective.
@@ 5 Related Work
Boltzmann generators are machine learning-based samplers that train likelihood-based models and employ inference-time SNIS to achieve consistent sampling of the target density (Noé et al., 2019). A major limitation of standard Boltzmann generators is the need to train the proposal model on a dataset of true density samples, motivating methods that transfer between molecular systems. Accordingly, Jing et al. (2022) propose a transferable Boltzmann generator operating on the torsion angles of small molecules, and Klein and Noe (2024) develop a Boltzmann generator operating on Cartesian coordinates that transfers between dipeptide systems. However, the scalability of such methods has remained limited due to the difficulty of designing expressive generative models that possess efficient and accurate likelihood evaluations. In particular the use of continuous normalizing flows implies a large cost to proposal likliehoods due to the need to integrate the vector field divergence (Grathwohl et al., 2019). Schopmans and Friederich (2025) replace the single-step SNIS of standard Boltzmann generators with a temperature-annealing sequence of normalizing flows, performing SNIS with each flow to sample from a given target distribution.
@@ Acknowledgments
Appendix
Appendix A Prose architecture Details
A.1 Tokenization and masking
[…]
with log-determinant of Jacobian given by
$\log\mathopen{}\mathclose{{\left|\frac{\partial f_{t}(z_{t})}{\partial z_{t}}}}\right|=-\sum_{i=1}^{N-1}\sum_{j=0}^{D-1}m[i]\cdot\alpha_{t}(z_{t}[:i])[i]_{j}.$ (12)
[…]
A.2 Permutations
All permutations are defined from the N-terminal to the C-terminal. In the residue-by-residue permutation the atoms are ordered such that each residue forms a contiguous sequence, with sidechain atoms immediately following the corresponding backbone atoms. In the backbone-first permutation the entire sequence of backbone atoms is placed at the start of the sequence before any sidechain atoms. Where constituent residues possess a branch or ring we also introduce a variant in which the branch ordering is flipped or ring traversal is inverted. The flip permutations are a simple inversion of the permutation, where padding tokens and not moved from their position at the end of the sequence. The specific sequence of permutations employed in Prose is presented in Table˜5.
Permutation Description
$\pi_{0}$ Backbone first
$\pi_{1}$ Residue–by–residue (flip)
$\pi_{2}$ Backbone–first (flip)
$\pi_{3}$ Residue–by–residue
$\pi_{4}$ Backbone–first (variant)
$\pi_{5}$ Residue–by–residue (variant, flip)
$\pi_{6}$ Backbone–first (variant, flip)
$\pi_{7}$ Residue–by–residue (variant)
A.3 Adaptive layer norm and transition
Appendix B Dataset
@@
Local energy minimization is performed with the Limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BGFS) algorithm. Energy minimization is followed by burn-in simulation of length $50\text{\,}\mathrm{ps}$, after which samples are collected every $1\text{\,}\mathrm{ps}$ (train) or $10\text{\,}\mathrm{ps}$ (test) until the simulation budget is exhausted. Full MD simulation parameters are provided in Table˜6.
@@
Appendix C Training configuration
All models are trained for $5\cdot 10^{5}$ iterations using a batch size of 512 with the AdamW optimizer [Loshchilov and Hutter, 2018]. We employ a cosine learning rate schedule in which the initial and final learning rates are a reduction of the maximal value by factor of 500, as well as exponential moving average with decay of 0.999. No overfitting was observed hence no early stopping was required. Given the large size of the training trajectories, we subsample to $10\text{\,}\mathrm{ps}$ per frame. Samples are normalized using values approximating the standard deviation of the 8AA data ($\sigma=0.35$), or if absent the 4AA data ($\sigma=0.28$), noting that a single value must be shared across systems of different dimensionality. An overview of all training configurations is provided in Table˜8.
Continuous Normalizing Flows
TarFlows
Following Zhai et al. [2024] and Tan et al. [2025] we use a learning rate of $1\cdot 10^{-4}$, weight decay of $4\cdot 10^{-4}$, and AdamW $\beta_{1},\beta_{2}$ of $(0.9,0.95)$. Data augmentation is applied as random rotations and Gaussian center of mass augmentation, in which every the entire system conformation is translated by a vector $c\sim\mathcal{N}(0,\sigma_{c}^{2}I_{3})$. The $\sigma^{2}$ value is chosen to match that of the prior, which has a center of mass variance $\sigma_{c}^{2}=\frac{1}{N}$ where $N$ is the number of atoms. Given $N$ is in our case variable for a single model trained on multiple systems, this augmentation is applied-per system before padding is applied. The architecture width and depth is provided in Table˜9.
Table 8: Overview of training configurations.
ECNF ECNF++ TarFlow / Prose
Learning Rate $5\cdot 10^{-4}$ $5\cdot 10^{-4}$ $1\cdot 10^{-4}$
Weight Decay 0.0 $1\cdot 10^{-2}$ $4\cdot 10^{-4}$
$\beta_{1},\beta_{2}$ 0.9, 0.999 0.9, 0.999 0.9, 0.95
EMA Decay 0.0 0.999 0.999
ECNF ECNF++ TarFlow Prose
Channels 128 256 384 384
Depth 9 9 8 8
Layers per block N/A N/A 8 8
Parameters (M) 1 4 115 285
C.1 Computational resources
ECNF++ TarFlow Prose
Training iterations / H100 hour 960 1132 260
Appendix D Importance sampling variants
D.1 Self-normalized importance sampling
Self-normalized importance sampling (SNIS) corresponds to the following estimator
@@ D.1 Self-normalized importance sampling
D.2 Continuous-time annealed importance sampling
@@
D.3 Discrete-time annealed importance sampling
@@
D.4 Sequential Monte Carlo
@@ E.1 Proposal sampling and likelihood evaluation
$\log p^{c}_{\theta}(x)=\log p_{\theta}(x)-\mathopen{}\mathclose{{\left[\log\mathopen{}\mathclose{{\left(\frac{\|c\|^{2}}{\sigma^{3}}}}\right)+\frac{\|c\|^{2}}{2\sigma^{2}}-\log\mathopen{}\mathclose{{\left(\sqrt{2}\Gamma\mathopen{}\mathclose{{\left(\frac{3}{2}}}\right)}}\right)}}\right]$ (37)
where $\Gamma$ is the gamma function. This adjustment seeks to account for the radial component introduced by translation non-equivariance. We additionally apply the same weight clipping threshold as in the ECNF when performing SNIS, or before SMC.
@@ E.2 Metrics
We compare generated samples to ground truth data, collected as defined in Appendix˜B, using empirical Wasserstein-2 distances. Given empirical distributions $\mu=\frac{1}{n}\sum_{i=1}^{n}\delta_{x_{i}}$ and $\nu=\frac{1}{m}\sum_{j=1}^{m}\delta_{y_{j}}$, the empirical Wasserstein-2 distance is defined as
@@
Following Raja et al. [2025], on the reference MD trajectory we run $k$-means ($k=20$) on the first two time-lagged independent components (TICs) fitted to the pairwise distances and dihedral angles of the peptide confirmations. Using these clusters, we obtain the occupancy distributions for the reference MD and the samples generated from Prose, and report their Jenson-Shannon divergence ($\mathrm{TICA}\text{‑}k\text{-}\mathrm{JSD}$), defined as
@@
where $P$ and $Q$ are discrete distributions, and $M=\frac{1}{2}(P+Q)$. Moreover, we repeat the same procedure using $k$-means only on the dihedral angles, and report this as $\mathcal{T}\text{‑}k\text{-}\mathrm{JSD}$. In both cases we fit the $k$-means clustering on the features (TICA projections or dihedrals) from full reference MD, but compute the metrics of generated samples against a subsampling of the reference trajectory as previously discussed. Note that JSD depends on the arbitrary choice of clustering (here $k=20$) and can be sensitive to binning resolution, whereas Wasserstein distances avoid discretization errors by operating directly on the continuous distributions.
@@ E.3 Sampling algorithm configurations
For both variants we generate a proposal set of $10^{4}$ samples. For the discrete variant, we perform 50 annealing steps, requiring two energy evaluations per step: one to update the samples and one for the Metropolis–Hastings correction. For the continuous variants, we perform 100 annealing steps. In both cases, resampling is performed at every step. For the continuous variant, Langevin dynamics is used with a step size $\sigma_{t}$ of $10^{-7}$ for dipeptides and tetrapeptides, and $\sigma_{t}$ of $10^{-8}$ for octopeptides. Details of the formulation are found in Section˜D.2. For the discrete variant, we apply Langevin dynamics with a step size of $10^{-5}$, followed by a Metropolis-Hastings step to accept or reject proposals. The step size is adaptively updated to maintain an acceptance rate of approximately 60%, under the assumption of sufficient smoothness in the intermediate densities. Further details of discrete-time AIS are found in Section˜D.3.
@@ E.4 Additional baseline configurations
The inference code of Lewis et al. [2024] was employed. BioEmu does not directly model all-atom resolution hence hpacker is employed to introduce the side chains before energy minimization. The codebase of [Lewis et al., 2024] was adapted to use the same Amber14 forcefield as ManyPeptidesMD, on which the $10^{4}$ energy evaluation budget was spent on per-sample minimization. We additionally experimented with equilibration to target the larger $10^{6}$ energy budget but could not achieve superior results to minimization-only hence this was not included. The adapted codebase for this baseline is provided at https://github.com/transferable-samplers/BioEmu.
@@ E.5 Computational resources
Appendix F Supplementary experimental results
F.1 Unweighted proposal performance
@@ F.2 Additional ablations
We consider three further ablations of the Prose architecture. Firstly, we replace the adaptive conditioning and transition blocks with (i) deeper transformation blocks in which the 8 transformer layers are increased to 20 (ii) a wider transformer block in which the dimension is increased from 384 to 576. In both cases the increased depth / width was defined to approximately match the parameter count of Prose. We additionally ablate the use of lookahead conditioning, in which the atom token $z[i]$ is conditioned not only on $A_{i},R_{i},P_{i},L_{i}]$ but also $A_{i+1},R_{i+1},P_{i+1},L_{i+1}]$, with conditioning information for this pair of indexes mixed using a small MLP. This was motivated by the observation that given naive conditioning the causal masking implies updates to $z[i]$ are computed without knowledge of $A_{i},R_{i},P_{i},L_{i}]$. Preliminary experiments up to tetrapeptides suggested look-ahead conditioning to be beneficial hence it was included in the final Prose model.
@@ F.3 Dataset scaling
We explore the effect of dataset size on the performance of Prose. The full ManyPeptidesMD dataset contains 21,700 sequences simulated for $200\text{\,}\mathrm{ns}$ each. We train a further four models wherein (i) the trajectories are limited to $50\text{\,}\mathrm{ns}$ and $100\text{\,}\mathrm{ns}$ respectively, and (ii) the set of sequences is reduced to 25% and 50%. For the 25% and 50% sequence reduction variants sequence reduction is applied non-linearly with a simple estimate of relative cost, to avoid excessively removing the relatively cheap simulations of short sequences. In Fig.˜9 we present metrics for these models. We note the $E\text{‑}\mathcal{W}_{2}$ to lack any meaningful trend, and the $\mathcal{T}\text{‑}\mathcal{W}_{2}$ to have limited sensitivity to the dataset variation, despite a slight trend towards larger data subsets improving the metrics. Most interesting of these plots is the $\mathrm{TICA}\text{‑}\mathcal{W}_{2}$ with a clear trend indicating the full $200\text{\,}\mathrm{n}\mathrm{s}$ trajectories to be beneficial, whilst using only 50% of the sequences is in fact superior to the full dataset. We caveat these results with the observation that Prose may indeed be operating in the compute-bound regime; even subsampled to $10\text{\,}\mathrm{ps}$ per frame the ManyPeptidesMD training dataset contains sufficient data for over $8\cdot 10^{5}$ training iterations (batch size 512) without repeating a single data sample, in excess of the training budget of $5\cdot 10^{5}$ permitted to the Prose models in this paper.
Figure 9: Wasserstein-distance metrics for Prose trained on variants of ManyPeptidesMD. Upper row: ManyPeptidesMD contains $200\text{\,}\mathrm{ns}$ trajectories for all training sequences, we train on a variant with $50\text{\,}\mathrm{ns}$ and $100\text{\,}\mathrm{ns}$ respectively. Lower: ManyPeptidesMD contains 21,700 sequences, we train using 25% and 50% of the total. Evaluation metrics computed using SNIS on 30 octopeptide sequences.
F.4 Sequence-length extrapolation
YQNPDGSQA (9AA) Chignolin (10AA)
UniSim $10^{5}$ 6.00 0.84 267.68 6.48 0.20
BioEmu 160.52 4.52 1.14 198.90 5.14 0.65
Prose Proposal $10^{9}$ 3.91 1.65 $10^{9}$ 3.63 0.96
Prose SNIS 0.0049 23.79 3.85 1.94 0.0001 832.59 4.35 1.25
Prose (self-improve) Proposal $10^{9}$ 3.73 1.95 $10^{9}$ 3.94 1.13
F.5 JSD evaluation